Can you help me with the following problem
A personnel manager is interviewing potential employees in order to fill two vacancies. The probability that the interviewee has the necessary qualities and accepts an offer is 0.8.
a) What is the probability that it will be necessary to interview exactly four people?
b) What is the probability that it will be necessary to interview less than four people?
Answer #1

a)

probability that it will be necessary to interview exactly four people

=P(Exactly one is suitable in first 3 and 2nd on 4th interview)=(3C1)*(0.8)^2*(0.2)^2=3*0.64*0.04=0.0768

b)

probability that it will be necessary to interview less than four people

=P(first 2 are selected)+P(Exactly one of first 2 and 3rd one is selected)

=0.8*0.8+(2C1)*0.8^2*0.2

=0.64+2*0.64*0.2

=0.8960